OpenStudy (anonymous):

The expressions 22x^3+15x^2-4x-6 and 2+14x-12x^2-5x^3 have the same remainder when each is divided by x-a. Show that 27a^3+27a^2-18a-8=0. Use the substitution y=3a to find the values of a.

OpenStudy (anonymous):

plug in x=a in both the polynomials and equate both .!

OpenStudy (anonymous):

u there ..? or i should elaborate it all ?

OpenStudy (anonymous):

Okay,sorry i replied so late.

OpenStudy (anonymous):

But,it's x-a not x=a

OpenStudy (anonymous):

i know ..what i am telling is the application of remainder theoremremainder the

OpenStudy (anonymous):

Okay,I will try it.

OpenStudy (anonymous):

yeah ..try if u get stuck ...m there .!

OpenStudy (anonymous):

What's next?

OpenStudy (anonymous):

\[22a^3+15a^2-4a-6=2+14a-12a^2-5a^3\]

OpenStudy (anonymous):

above i have simply put "a" in place of x and equated them then just simply by transposing like terms...u will get ..! ...\[27a^327a^2-18a-8=0 \]

OpenStudy (anonymous):

Okay Thanks